OpenStudy (anonymous):

What is the solution set of x2 + y2 = 26 and x − y = 6? {(5, -1), (-5, 1)} {(1, 5), (5, 1)} {(-1, 5), (1, -5)} {(5, -1), (1, -5)}

hartnn (hartnn):

plug in x = y+6 in x^2+y^2 =26 what do u get ?

OpenStudy (anonymous):

I don't know?

OpenStudy (shamil98):

If x = y + 6 and the equation is \[x^2 + y^2 =26\]

OpenStudy (shamil98):

you substitute what x is equal to in the other equation, and solve for y.

OpenStudy (shamil98):

\[(y+6)^2 + y^2 = 26\] Solve for y from there.

OpenStudy (anonymous):

but I don't know how to do that?

hartnn (hartnn):

do you know thee formula for \((a+b)^2\) ??

OpenStudy (anonymous):

i need help with trying to figure out how to solve it

OpenStudy (anonymous):

no i dont

hartnn (hartnn):

ohhh, sorry then i cannot help you. good luck!

jimthompson5910 (jim_thompson5910):

solve the second equation for x x - y = 6 x = 6 + y x = y + 6 ---------------- then plug this into the first equation x^2+y^2 =26 x^2+y^2 =26 (y+6)^2+y^2 =26 y^2 + 12y + 36 + y^2 = 26 y^2 + 12y + 36 + y^2 - 26 = 0 2y^2 + 12y + 10 = 0 2(y^2 + 6y + 5) = 0 y^2 + 6y + 5 = 0 (y+1)(y+5) = 0 I'll let you finish up